ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 68 Belabas, Karim, Université Paris-Sud, France PARI / GP The PARI / GP package is aimed at arithmeticians but should be of interest to mathematicians whose primary needs are speed (part of the system is written in assembly language) and ease of use. All standard number theoretical functions are included as well as a huge algebraic number theory module, with particular emphasis on explicit class field theory. The package is not primarily a computer algebra or symbolic manipulation engine. PARI is a C library which can be used directly for greater speed, or interactively through an interpreter (called GP), using mathematical concepts rather than computer language. The package runs on all major architectures and operating systems, it includes complete source code and is free for non commercial purposes. You can download PARI from http://pari.home.ml.org